OpenStudy (anonymous):

write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4i and square root of 3

OpenStudy (anonymous):

the other two zeros must be \[-4i,-\sqrt3\] multiply \[(x-4i)(x+4i)(x-\sqrt3)(x+\sqrt3)\]

OpenStudy (anonymous):

not nearly as hard as it looks \[(x+4i)(x-4i)=x^2+16\] and \[(x-\sqrt3)(x+\sqrt3)=x^2-3\]

OpenStudy (anonymous):

your real job then is only to multiply \[(x^2+16)(x^2-3)\]

OpenStudy (anonymous):

So, it would end up being x^4 + 13x^2 - 48

OpenStudy (anonymous):

thank you! @satellite73

OpenStudy (anonymous):

yw